What is the x value of the solution to this system of equations? 3x+2y=6 2x+y=2
Answer by macston(5194)

.
2x+y=2
y=2-2x . Use this to substitute for y.
.
3x+2y=6 . Substitute for y.
3x+2(2-2x)=6
3x+4-4x=6
-x=2
x=-2
ANSWER: The x value is -2.
.
CHECK:
y=2-2x
y=2-2(-2)
y=2-(-4)
y=6
.
3x+2y=6
3(-2)+2(6)=6
-6+12=6
6=6
.
2x+y=2
2(-2)+6=2
-4+6=2
2=2